Dr. Owino needs two students to help him with a science demonstration for his class of 18 girls and 12 boys. He randomly chooses one student who comes in front of classroom. He then chooses a second student from those still seated. What is the probability that both students chosen are girls?Β 


1
Expert's answer
2021-07-06T17:10:32-0400

"P = \frac{18}{30} \times \frac{17}{29} \\

= \frac{51}{145} \\

= 0.352"